### v1

# Time dependence of neutrino quantum kinetics in a core-collapse supernova

Abstract
Our understanding of neutrino flavor conversion in the supernova core is still preliminary, despite its likely relevance to the neutrino-driven supernova mechanism. We present multi-angle and multi-energy numerical simulations of neutrino quantum kinetics within a spherically symmetric shell in the proximity of the region of neutrino decoupling. We rely on inputs from a one-dimensional core-collapse supernova model with a mass of and find that, at early post-bounce times ( s), no crossing is present in the angular distribution of the electron neutrino lepton number and flavor conversion develops due to vacuum mixing. Angular crossings appear for s and fast flavor conversion leads to flavor equipartition, with the spectral energy distribution of () and () becoming comparable. Notably, flavor equipartition is not a generic outcome of fast flavor conversion, but rather a consequence of the relatively similar properties of neutrinos of different flavors characterizing the late accretion phase. Artificially tweaking the collision term to introduce an electron lepton number angular crossing for s, we observe that flavor equipartition is not achieved. While our findings are restricted to a specific supernova model, they suggest a rich phenomenology of flavor conversion in the supernova core as a function of the post-bounce time which needs to be further explored to assess its impact on the explosion mechanism.
1 Introduction
In neutrino dense astrophysical sources, flavor evolution is significantly affected by the coherent forward scattering (refraction) of neutrinos with each other [1, 2]. Neutrino self-interaction, however, makes the neutrino equations of motion nonlinear and the consequent flavor evolution displays a rich phenomenology that we are still far from understanding [3, 4, 5, 6, 7].
The modification of neutrino flavor evolution due to refraction with other neutrinos can be broadly classified into slow [8, 9, 4], fast [10, 11, 12, 13, 14, 6, 7], and collisional [15]. Slow flavor evolution is differentiated from the fast one by the role of the vacuum frequency. In fact, fast flavor evolution can manifest in the limit of vanishing vacuum frequency and it is triggered by a crossing in the electron lepton number (ELN) angular distribution of neutrinos [14, 13, 16], whereas slow flavor evolution requires a nonzero vacuum term [9, 17]. Fast flavor evolution, as the name suggests, occurs over very small time scales—the characteristic frequency being directly proportional to the number density of neutrinos. On the other hand, collisional instabilities should be triggered by neutrino collisions with the medium in the high-density region in the proximity of neutrino decoupling [18, 19, 20, 21], but their relevance on flavor conversion physics within realistic astrophysical environments remains to be assessed [22, 23].
Neutrinos play a crucial role in the core-collapse supernova mechanism [24, 25], as recently confirmed by multi-dimensional hydrodynamic simulations [26, 27, 28, 29]. Whether neutrino flavor conversion, triggered by any of the flavor instabilities highlighted above, affects the core collapse mechanism is the subject of ongoing research [30, 31, 32]. Preliminary work suggests that, if flavor equilibration should occur as a result of such instabilities, this would facilitate or hinder the supernova explosion according to the mass of the collapsing star [31]. Because of the technical challenges linked to the solution of the neutrino quantum kinetic equations (QKEs) and the vast difference in the characteristic scales typical of the source hydrodynamics and the ones entering the neutrino equations of motion, a self-consistent solution of this problem is not yet available [6, 7]. Nevertheless, it is possible to search for general trends by solving the QKEs within a simulation shell, taking into account the interplay among flavor evolution, collisions, and advection. In this case, a quasi-steady state configuration is found in the aftermath of flavor conversion [33, 34, 35, 32, 22, 36, 37, 38].
An open question concerns whether flavor equipartition is a general solution of the QKEs. While recent work seems to point in this direction [39, 36, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], the boundary conditions as well as the ELN distribution entering the solution of the QKEs could drastically affect the final flavor configuration [37, 33, 34]. In this paper, we investigate the neutrino flavor evolution in spherical symmetry, relying on input from several post-bounce time snapshots within s of a one-dimensional hydrodynamic simulation of a core-collapse supernova with a mass of [50]. We find that, in the early accretion phase ( s), flavor conversion is driven by vacuum mixing; while, in the late accretion phase ( s), equipartition between the neutrino spectral energy distributions of and as well as and . Yet, we confirm early findings [33, 34, 37] showing that flavor equipartition is not a general flavor outcome for any given unstable flavor configuration.
This paper is organized as follows. Section 2 introduces the neutrino QKEs and outlines the approach adopted to model the collision term as well as the simulation setup. In Sec. 3, we describe the main features of our benchmark one-dimensional supernova model, compute the neutrino angular distributions of all flavors employing a multi-energy transport scheme and illustrate as they evolve as functions of the post-bounce time. Our main findings are presented in Sec. 4. Finally, we discuss our results in Sec. 5 and conclude in Sec. 6.
2 Neutrino quantum kinetics in spherical symmetry
In this section, we introduce the neutrino equations of motion. We also outline the numerical techniques adopted to solve the neutrino QKEs.
2.1 Quantum kinetic equations
For simplicity, we consider two neutrino flavors ( and with ), and the corresponding antineutrinos ( and ). We model our neutrino ensemble in terms of density matrices for neutrinos and antineutrinos for each radial location (), polar angle (), and energy (), as a function of time () and under the assumption of spherical symmetry, see Fig. 1 111Note that the polar angle is defined with respect to the radial direction at a given point and it is not the emission angle used in the early papers on neutrino-bulb model [8, 9, 4].. The diagonal elements of the density matrices ( with and ) represent the neutrino occupation number and are normalized such gives the local neutrino number density of flavor at a given radius ().
The equations of motion that describe the evolution of the (anti)neutrino density matrices are:
The total derivative on the left hand side of Eqs. 2.1 and 2.2 consists of a partial derivative with respect to time and the advective term, which is in spherical symmetry
The advective term is such that, in the absence of any term other than the total derivative, neutrinos travel along a straight line. General relativistic effects are neglected in Eqs. 2.1 and 2.2, qualitatively they should not affect the interplay between neutrino flavor conversion and transport effects on small scales [32].
On the right hand side of Eqs. 2.1 and 2.2, the commutator of the Hamiltonian and the density matrix takes into account neutrino flavor conversion physics. The Hamiltonian consists of three components: the vacuum, matter, and self-interaction terms. The matter term effectively suppresses the mixing angle in the vacuum term; hence, we choose to adopt an effective vacuum mixing angle, ignoring the matter term:
The vacuum and self-interaction terms are defined as
The vacuum frequency is with eV2, and rad denotes the matter suppressed mixing angle. The self-interaction Hamiltonian, unlike the vacuum Hamiltonian, depends on and is obtained integrating over all momentum modes. The integration over the azimuthal angle results in a constant factor that has been absorbed in the definition of the density matrices. On the other hand, the self-interaction Hamiltonian does not depend on energy. Following the attenuation method introduced in Ref. [51], we rescale the strength of neutrino self-interaction by a constant factor . The choice of is such that the self-interaction strength is always larger than at neutrino decoupling and the quasi-steady state flavor state configuration is likely not affected, but the time needed for reaching such a configuration changes with the computational time being reduced [52, 53].
The collision term on the right hand side of Eqs. 2.1 and 2.2, represented by and , includes three contributions: the emission term (), the absorption term (), and the direction changing term () [54, 55, 28]. For all flavors of neutrinos, pair production and Bremsstrahlung contribute to the emission and absorption terms, whereas the beta processes only contribute to electron-type neutrinos and antineutrinos. The direction changing term is given by the neutral current interaction between neutrinos and nucleons, which are assumed to be elastic. The inelastic contribution is negligible for the electron type neutrinos and thus does not affect the presence of ELN crossings. The collision terms are defined as follows:
we use the symbol to denote the elementwise multiplication:
The emission and absorption terms are related by Kirchoff’s law for each energy, , where is the Fermi-Dirac energy distribution. An analogous expression holds for . The terms and are included out of completeness and take into account the anisotropic nature of direction changing neutral current interactions. The collision terms used in this paper are the multi-energy extension of the ones presented in Appendix A of Ref. [23], which in turn are based on Refs. [56, 57].
2.2 Simulation setup
In order to solve the QKEs, we use the Julia implementation of the Adam-Moultan multistep integrator, with adaptive order and adaptive step size for the time variable 222The length scale characteristic of our neutrino ensemble is not given by the inverse of the self-interaction potential, rather it is a combination of the collision term and the self-interaction potential [33, 34]. Yet, it is important to ensure that the time step is much smaller than the inverse of the neutrino self-interaction strength; we do this by implementing an adaptive step size. (VCABM solver). The spatial derivatives in the advective term are calculated using the central difference method. To discretize the density matrix, we employ angle bins, energy bins for MeV, and radial bins for (with ans being the innermost and outermost radii of our simulation shell; see Fig. 1).
Our goal is to compute the “quasi-steady state” configuration achieved by the system 333A quasi-steady state configuration does not coincide with the solution of the boundary value problem obtained by setting the time derivative to zero. The quasi-steady state is also not approximately close to the solution of that boundary value problem. The reason for this counter-intuitive assertion is that, if Eqs. 2.1 and 2.2 are evolved for a long enough time, , and reach a steady state, but the phase of does not converge towards a steady state. This explains why we need to evolve the system as a function of time instead of attempting to find a solution to Eqs. 2.1 and 2.2 setting the time derivative to zero.. The latter should be independent of the initial configuration, but the choice of the initial configuration does affect the efficiency and numerical stability of the numerical solution. As pointed out in Refs. [33, 34], the most efficient way to tackle this issue numerically is to first solve Eqs. 2.1 and 2.2 for to obtain the “classical steady state” solution (i.e., we solve the classical Boltzmann equation). We then use the classical steady state solution as the initial configuration to solve Eqs. 2.1 and 2.2 for and . This approach has two advantages. First, it is easier to highlight the impact of flavor evolution; second, the initial classical steady state configuration is not too different from the final quasi-steady state. We stress, however, that the choice to use the classical solution as the initial configuration is not unique, in principle it might be possible to devise a better initial configuration.
For obtaining the classical steady-state solution as well as the quasi-steady state one, the radial range has to be chosen such that the neutrino spectral energy distributions coincide with Fermi-Dirac distributions at the minimum radius (); whereas the maximum radius () is chosen such that the classical solution has negligible backward flux [23], see also Fig. 1 and Table 1. These conditions serve as boundary conditions for Eqs. 2.1 and 2.2.
(s) | (km) | (km) |
25 | 150 | |
25 | 100 | |
22 | 57 | |
20 | 35 | |
17 | 32 | |
16 | 31 |
The hydrodynamic and thermodynamic supernova properties employed to compute the classical and quasi-steady state configurations are extracted from a one-dimensional core-collapse supernova simulation whose features are illustrated in Sec. 3.
3 Benchmark core-collapse supernova model
We rely on the outputs of a one-dimensional hydrodynamic simulation of a core-collapse supernova, with gravitational mass and SFHo nuclear equation of state without muons [50]. In this model, the effects of proto-neutron star convection have been taken into account through a mixing-length approximation [3, 58].
We focus on few selected post-bounce times (, , , , , and s) and rely on static hydrodynamic backgrounds and thermodynamical quantities for each time snapshot to compute the classical steady-state neutrino configuration. Figure 2 shows the radial evolution of the main supernova properties entering the collision term in Eqs 2.1 and 2.2.
Since the neutrino angular distributions are not provided as an output of our benchmark hydrodynamic simulation, we compute them following the same procedure illustrated in Sec. III of Ref. [23], i.e. we set and as indicated in Table 1 for each snapshot and solve Eqs. 2.1 and 2.2 for . In contrast to Ref. [23], we compute the angular distributions of (anti)neutrinos solving Eqs. 2.1 and 2.2 for multiple energy modes (see Sec. 2.2).
The left panels of Figs. 3 and 4 display the contour plot of the energy-integrated in the plane spanned by and in the absence of flavor conversion (i.e., the classical steady state configuration). One can see that, at small radii (before neutrino decoupling), the angular distributions of neutrinos and antineutrinos are isotropic with an overall excess of . At larger radii, (anti)neutrinos start to decouple from matter entering the free streaming regime. As expected, because of the overall decrease of baryon density, neutrinos decouple at smaller radii as the post-bounce time increases. As (anti)neutrinos approach the free-streaming regime, an ELN crossing forms for s (visible from the white contour separating the red from the blue regions in the left panels of Fig. 4). The presence of an ELN crossing is a necessary condition for flavor instability and becomes a sufficient condition, if periodic boundary conditions are assumed [14, 59].
Figure 5 shows the energy-integrated ELN angular distributions, extracted at , as functions of for our selection of post-bounce time snapshots. We can see that ELN crossings appear for s and are not present at early times. However, there is no monotonic trend as a function of radius for what concerns the ELN appearance at different . It is worth noticing that Ref. [23] found ELN crossings for all considered post-bounce time snapshots because it considered a single-energy solution of Eqs. 2.1 and 2.2 for the sake of simplicity. We use this set of classical steady state configurations to investigate the differences in the flavor conversion outcome, as illustrated in the next section.
4 Time dependent neutrino quantum kinetics
In this section, we present our results on the solution of the QKEs for different supernova time snapshots. First, we focus on the early accretion phase ( s) and then we investigate the quasi-steady state configuration during the late accretion phase ( s). Finally, we artificially modify the collision term to investigate whether flavor equipartition is a generic outcome of flavor conversion.
4.1 Early accretion phase: no ELN crossings, minimal flavor conversion
Although we do not find ELN crossings at early times ( s), this does not imply the absence of flavor evolution as displayed in the right panels of Fig. 3 for and s. Such flavor evolution is triggered by the vacuum mass term. Earlier work concluded that slow flavor evolution should be suppressed at early post-bounce times due to the large ratio of the self-interaction strength and the vacuum term [60, 61, 62, 63, 64, 65]. However, as recently shown in Ref. [66], if the ELN distribution is on the verge of developing a crossing, the vacuum term can induce a flavor instability even when the self-interaction strength is large (inducing a crossing in the effective ELN distribution). This is exactly what happens at and s. We verify this conjecture by performing the linear stability analysis without rescaling the self-interaction strength. As shown in Fig. 6, a non-zero growth rate of the off-diagonal term of the density matrix is found.
The right panels of Fig. 3 show that, unlike for the flavor instability induced by an ELN crossing (see Sec. 4.2), flavor equipartition does not take place. This can be attributed to a growth rate that does not dominate over the advective length scale as in the case of fast flavor instability and the different shape of the ELN angular distribution (which exhibits a larger difference between the properties of and with respect to later post-bounce times).
4.2 Late accretion phase: appearance of ELN crossings and fast flavor conversion
The right panels of Fig. 4 show the contour plots of the energy-integrated at the simulation time s in the plane spanned by and and in the presence of flavor conversion for , and s (from top to bottom respectively). We find that flavor conversion develops in the proximity of the ELN crossings (the latter being denoted by the white region in between the red and blue ones) for , and s. Flavor conversion is therefore driven by fast flavor instabilities, as expected from the preliminary findings of Ref. [23] in the single-energy approximation. In fact, Ref. [23] found that the growth rate of the fast flavor instability is four orders of magnitude larger than the collisional instability one for these time snapshots (cf. their Fig. 6). Reference [23] also found that collisional instabilities are present within small spatial regions at smaller radii, yet we do not see any significant flavor conversion linked to them.
In order to investigate whether flavor equipartition is achieved as a result of flavor conversion, Fig. 7 displays the energy-integrated for the quasi-steady-state solution obtained after flavor conversion for neutrinos (antineutrinos) on the left (right) panels. Note that, in the region of neutrino trapping, there is an excess of over due to the positive chemical potential of ; while the opposite is true for antineutrinos. For all analyzed configurations for s, we find flavor equipartition is achieved for neutrinos and antineutrinos. This is also visible from the left panels of Fig. 9, where the radial evolution of the neutrino number density of the different flavors after flavor conversion (quantum solution) is compared to the one obtained neglecting flavor conversion (classical solution). In particular, as shown in the right panels of Fig. 9 for (top two panels) and s (bottom two panels), the angle-integrated energy distribution of () is identical to the one of () in the quasi-steady state configuration and at large radii ( has been chosen to extract the spectral distributions plotted in Fig. 9, although any radius after neutrino decoupling shows a similar trend).
In the left panels of Fig. 9, the decoupling radii for each neutrino flavor are marked by vertical lines. The decoupling radius has been computed as the radius at which the flux factor is equal to [33, 67, 68]. We can see that the decoupling radius is minimally affected by flavor conversion (cf. dashed vs. solid lines); in fact, as shown in Ref. [33], the impact of flavor conversion on the decoupling physics strongly depends on the initial ELN distribution and how similar the electron and non-electron flavor emission properties are in the absence of flavor conversion.
4.3 Flavor equipartition is not a generic flavor outcome of fast flavor conversion
The neutrino properties obtained in the late accretion phase tend to be similar to each other for the different post-bounce times that we have considered (cf. Figs. 3, 4 and 9). In order to assess whether flavor equipartition should be expected to be the general flavor outcome of any quasi-steady state configuration, we increase the emission term of at s by to induce an ELN crossing. Solving the QKEs numerically, we find that flavor equipartition is not reached in the quasi-steady state solution, as shown in Fig. 10.
This flavor outcome can be explained by looking at the number density of and in the region of flavor instability: the two differ substantially compared to the configurations obtained at later post-bounce times for which flavor equipartition was found. The observed quasi-steady state configuration is due to the fact that the collision term tends to enhance the number density and suppress the one. On the other hand, flavor conversion tends to push and towards each other. As a result, the number density tends to change more (moving upwards) in the quasi-steady state configuration than the one with respect to the classical solution, being the impact of the collision term smaller for the non-electron flavors. As a consequence, the decoupling radii are also more largely affected by flavor conversion than observed for the other flavor configurations (cf. Fig. 4).
Figure 11 shows the spectral energy distributions extracted for selected time steps at km and km, in order to investigate the temporal evolution of the quasi-steady flavor configuration. The neutrino spectra continue to oscillate in time, changing by as much as . However, the system is approximately in partial equipartition for the neutrino sector and not for the antineutrino sector at km (cf. left panel). This trend changes as a function of radius, e.g. the opposite is true for km (cf. right panel), and in general clearly shows that flavor equipartition is not achieved for this flavor configuration.
We note that the time over which flavor evolution occurs is much shorter than the time scale over which the collision term changes. The continued temporal evolution of the spectral energy distributions suggests that the physically relevant quantities with potential feedback on the source physics should be the average of the bands in Fig. 11.
5 Discussion
In order to assess the possible impact of our findings on neutrino flavor conversion physics on the supernova mechanism, Fig. 12 shows the ratio of the neutrino heating rates obtained relying on the quasi-steady state configuration when the QKEs are solved with and without the commutator term encapsulating the flavor conversion physics. The heating rate is defined as
where denotes the energy or or , MeV denotes the -value of the -reaction, MeV is the mass of the electron, is the characteristic neutrino interaction cross section ( cm2), is the electron Pauli-blocking factor, and is the axial coupling.
Figure 12 shows that, for all cases except s (for which we do not see any flavor conversion), an increase in the heating rate by – occurs. This is not surprising since the flavor evolution increases the number density of electron type neutrinos in the high energy tail of the spectral distribution, and a larger heating rate is expected for larger neutrino energy. While we do not compute the neutrino cooling rate and do not take into account the feedback that the neutrino flavor conversion physics could have on the supernova hydrodynamic background, our results suggest a non-trivial impact of neutrino flavor conversion physics on the supernova mechanism, supporting earlier parametric studies [31, 30, 32].
Our findings should be considered with caution as they have been obtained relying on a spherically symmetric core-collapse supernova model. While we observe a non-trivial appearance of ELN crossings as a function of the post-bounce time in the classical solution of the Boltzmann equations, multi-dimensional core-collapse supernova simulations show evidence of large scale asymmetries in the neutrino emission (e.g. due to the LESA instability [69]) and the location of flavor instabilities may be impacted [70].
In order to facilitate the numerical solution of the QKEs, we choose to attenuate the self-interaction strength by a constant factor (see Eq. 2.6), nevertheless for all pot-bounce times we have that at neutrino decoupling. When flavor equipartition is reached, it is unlikely that the rescaling of the self-interaction strength can affect the results. However, for the earlier accretion phase ( s), flavor equipartition is not reached, and it is possible that this scaling might slightly affect the exact final flavor outcome and the time needed to reach a quasi-steady state solution.
The energy-dependent solution of the QKEs has consequences on the development of ELN crossings (cf. the corresponding single-energy solution presented in Ref. [23]) and therefore the presence of flavor instabilities; yet, for all post-bounce times, we find that a quasi-steady state configuration is achieved—cf. also Refs. [71, 51]. These findings support recent attempts aiming at forecasting the quasi-steady state flavor configuration due to flavor conversion or designing methods to incorporate such physics into neutrino-radiation-hydrodynamic simulations [41, 42, 36, 72, 35, 73, 74]; on the other hand, caution should be adopted in the forecast of the quasi-steady state neutrino configuration because of the dependence of the latter on the boundary conditions of the simulation box and features of the ELN distribution [37, 41, 34].
We find that flavor equipartition is obtained for s. Such results may seem to confirm recent work in this direction [39, 36, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. However, as shown through the flavor configuration obtained by tweaking the collision term in Sec. 4.3, we stress that flavor equipartition depends on the initial conditions and should not be considered a general outcome of global flavor simulations [37, 33, 34].
Similar to Ref. [22], we explore the flavor conversion physics for a range of neutrino flavor configurations linked to different post-bounce times. While our results are in general agreement with the ones of Ref. [22], we scan a broader range of post-bounce times highlighting time-dependent features of the flavor outcome. Moreover, differently from Ref. [22], we do not artificially tweak the electron abundance to favor the appearance of ELN crossings; this has implications on the relative relevance between collisional and fast flavor instabilities. We also confirm the preliminary findings of Refs. [23, 70, 22], where it was concluded that collisional instabilities do not affect the flavor composition significantly.
6 Conclusions
Understanding the neutrino flavor distribution resulting from flavor conversion occurring in the dense core of a supernova is crucial in order to investigate the impact of this physics on the supernova explosion mechanism. We numerically solve the neutrino quantum kinetic equations (QKEs) in a multi-energy and multi-angle framework and within a spherically symmetric shell, tracking neutrino decoupling.
We solve the QKEs adopting as inputs static hydrodynamic backgrounds extracted from a one-dimensional core-collapse supernova simulation with a mass of at post-bounce times , , , , , and s. First, neglecting flavor conversion, we compute the angular distributions of neutrinos and antineutrinos, following Ref. [23]. Differently from Ref. [23] and because of the multi-energy numerical solution of the QKEs, we find that crossings in the electron neutrino lepton number (ELN) angular distributions appear for s, but not at earlier post-bounce times.
We adopt the classical steady-state solution to solve the QKEs, this time taking into account neutrino flavor conversion. Fast flavor conversion is responsible for making the spectral energy distribution of () equal to the one of () for s, i.e. flavor equipartition is reached. Despite the changes in the neutrino energy and angular distributions of all flavors, we find that the flavor-dependent neutrino decoupling radii are negligibly affected for the investigated flavor configurations; this is because of the relatively similar properties between the electron and the non-electron flavors characterizing the late accretion phase that facilitate flavor equipartition, but overall negligibly affect the energy and angle integrated flavor-dependent number density at decoupling. On the other hand, ELN crossings are not found for s, but flavor conversion develops because of vacuum mixing, which effectively induces a crossing in the ELN angular distributions.
Remarkably, flavor equipartition is not a generic outcome of fast flavor conversion, but rather the result of the interplay between collisions, flavor conversion and advection, when the neutrino emission properties of the electron and non-electron flavors are reasonably comparable to each other. Modifying the collision term artificially for a time snapshot extracted at s to induce a crossing in the ELN angular distribution, we find that flavor equipartition is not achieved. This is because the differences in the emission properties of different flavors are larger in the early accretion phase and the interplay between flavor conversion and collisions cannot equilibrate the electron and non-electron flavors.
Our findings should be expanded to draw conclusions on the flavor conversion phenomenology expected in a broader range of core-collapse supernova models. Moreover, a self-consistent implementation of flavor conversion in supernova hydrodynamic simulations may dynamically change the neutrino emission properties as a function of the post-bounce time, in turn possibly affecting the flavor conversion phenomenology. Despite its intrinsic limitations, our work highlights a very interesting, non-trivial flavor phenomenology in core-collapse supernovae, calling for further investigation in order to assess its impact on the neutrino-driven explosion mechanism and nucleosynthesis.
Acknowledgments
This project has received support from the Danmarks Frie Forskningsfond (Project No. 8049-00038B), the European Union (ERC, ANET, Project No. 101087058), and the Deutsche Forschungsgemeinschaft through Sonderforschungbereich SFB 1258 “Neutrinos and Dark Matter in Astro- and Particle Physics” (NDM). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.